hartnn (hartnn):

\(\Large \dfrac{π^4}{90}=\dfrac{1}{1^4} +\dfrac{1}{2^4} +\dfrac{1}{3^4} +\dfrac{1}{4^4} +⋯\)

OpenStudy (zzr0ck3r):

pretty

hartnn (hartnn):

I have proved that \(((π-x )/2)^2=π^2/12+[1/1^2 cos 1πx+1/2^2 cos 2πx+⋯]\) from which i could deduce that \([1/1^2 +1/2^2 +1/3^2 …]=(-π^2)/6 \) now the 2nd part of deduction asks me to prove that.

hartnn (hartnn):

thats supposed to be \(\Large +\pi^2/6 \)

hartnn (hartnn):

what could i do for the 2nd deduction ?

OpenStudy (kirbykirby):

Are you familiar with \[ \sin(\pi z)=\pi z\prod_{n=1}^{\infty}\left(1-\frac{z^2}{n^2} \right), z\in \mathbb{C}\]

hartnn (hartnn):

yes, but i need to deduce that by using series for (pi - x)^2/2 its the 2nd part of that question. i cannot use series for sin pi z

OpenStudy (sidsiddhartha):

u can use parseval's indentity

hartnn (hartnn):

question : find fourier expansion for [(pi-x)/2 ]^2 and deduce those 2

OpenStudy (sidsiddhartha):

there's a formula \[\int\limits_{-c}^{c}[f(x)^2]dx=c[\frac{ a_0^2 }{2 }+\sum_{n=1}^{\infty}(a_n^2+b_n^2)]\]

hartnn (hartnn):

yeah, i saw that, let me try that and get back to you

OpenStudy (sidsiddhartha):

x --0 to (pi) ??

hartnn (hartnn):

0 to 2pi

hartnn (hartnn):

\(\Large ((\pi-x)/2 )^2 = \pi^2/12 +\sum \limits _{n=1}^\infty (1/n^2) \cos n\pi x \) then what would be next ? how would i apply parseval here ?

hartnn (hartnn):

bn = 0 an = 1/n^2

hartnn (hartnn):

\(\Large \int\limits_{0}^{2\pi }[(\dfrac{(\pi-x)^2}{2})^2]dx=2\pi [\frac{ (\pi^2 /6)^2 }{2 }+\sum_{n=1}^{\infty}(1/n^4 )]\) like this ?

hartnn (hartnn):

and i need to solve the integral on left side ?

hartnn (hartnn):

too much work for a couple of marks

OpenStudy (sidsiddhartha):

yes and it will be a dangerous one :(

OpenStudy (sidsiddhartha):

hey what is a_0=pi^2/6 ?

hartnn (hartnn):

yes

hartnn (hartnn):

whats the exact formula for parsevals identity which i can use here...i would work it out, just wanna know the formula that i can use

OpenStudy (sidsiddhartha):

the formula i wrote that's the exact form

hartnn (hartnn):

ok, got my mistake left side = pi^5/40

hartnn (hartnn):

still i get -pi^4/720 ...

OpenStudy (sidsiddhartha):

yes left side pi^5/40

hartnn (hartnn):

then i get i get -pi^4/720

OpenStudy (sidsiddhartha):

yes i'm also getting the same thing:(

hartnn (hartnn):

that formula is for -c to c is there some different formula for 0 to a ?

OpenStudy (sidsiddhartha):

there could be three cases i think 1) if 0<x<2c \[\int\limits_{0}^{2c}[f(x)]^2dx=c[\frac{ a_0^2 }{ 2}+\sum_{}^{}a_n^2+b_n^2]\] 2)if 0<x<c \[\int\limits_{0}^{c}[f(x)]^2dx=c/2[\frac{ a_0^2 }{ 2 }+\sum_{}^{}a_n^2]\] this is like halfrange cosine series 3)0<x<c (half range sine) \[\int\limits_{0}^{c}[f(x)]^2dx=c/2[\frac{ a_0^2 }{ 2 }+\sum_{}^{}b_n^2]\]

hartnn (hartnn):

so i just try using pi instead of 2pi on right...? bn = 0 so 3) case =0

hartnn (hartnn):

yep, just using pi gave me exact pi^4/90

OpenStudy (sidsiddhartha):

yeah atlast :)

hartnn (hartnn):

seriously, too much work for a couple of marks

OpenStudy (sidsiddhartha):

yeah :) are u preparing for ies or something

hartnn (hartnn):

no no... i am writing past paper solutions ..

OpenStudy (sidsiddhartha):

past papers of?

hartnn (hartnn):

sem 3 maths, mumbai university

OpenStudy (sidsiddhartha):

awesome :3